OpenStudy (anonymous):

Write the equation of the line, in standard form, that has an x-intercept of 2 and is parallel to 2x + y = -5. In

OpenStudy (displayerror):

You're told that the x-intercept is 2. What is the coordinate at that point? You're also given the equation of a line, \(2x + y = -5\). What is the slope of that line?

OpenStudy (anonymous):

m=-2

OpenStudy (displayerror):

Right! The coordinate of the x-intercept is?

OpenStudy (anonymous):

I'm not sure

OpenStudy (displayerror):

The x-intercept is where the line crosses the x axis. Likewise, the y-intercept is where the line crosses the y-axis. What is the y-value at any point along the x-axis? (Remember that the x-axis is the horizontal axis)

OpenStudy (anonymous):

Using the given equation y=0

OpenStudy (displayerror):

Right! But that's true regardless of what equation you're using. Remember that the point where the x-axis and y-axis intersect/cross is \((0,0)\). Starting from that point \((0,0)\), if we go or down along the y-axis, we'll have points given by \((0,y)\); if we go left or right along the x-axis, we'll have points given by \((x,0)\). Back to the question: If we're told that the x-intercept is 2, what is the coordinate/point that represents that?

OpenStudy (anonymous):

2,0

OpenStudy (displayerror):

Great! Now what is the relationship between the slope of two parallel lines (which is the case for this problem)?

OpenStudy (anonymous):

What do you mean by that?

OpenStudy (displayerror):

Remember that for two lines that are perpendicular, their slopes are negative reciprocals of each other (for example, \(m=-2\) and \(m=\dfrac{1}{2}\) and thus, multiply together to equal -1. What is the relationship between the slopes of two parallel lines?

OpenStudy (anonymous):

Same

OpenStudy (anonymous):

As in parallel have same slope

OpenStudy (displayerror):

Yep! Parallel lines have the same slope; perpendicular lines have slopes that are negative reciprocals of one another. Now you have all of the information that you need--a point \((2,0)\) and a slope \(m=-2\) (which we know because we have the slope of the equation that's given to you),. Now you can plug this information into the equation of a line to get your equation, then convert it into standard form \(Ax+By=C\)

OpenStudy (anonymous):

This is where I trip up

OpenStudy (displayerror):

There are several forms of the equation of a line: Slope-intercept form: \(y=mx+b\) >>>Use this when you have a slope and y-intercept available Point-slope form: \(y-y_1=m(x-x_1)\) >>>>Use this when you have a point \((x,y)\) and a slope Standard form: \(Ax+By=C\) >>>Usually you'll use one of the two forms above to get to the standard form. In this problem, the only information that we have is a point \((2,0)\) and a slope \(m=-2\). Which form of the equation for a line should we use?